I need help finding the vertex of the parabola... I'm just not sure how to break this equation down 3x^2+30x+71
Answer by Alan3354(69443)

I need help finding the vertex of the parabola... I'm just not sure how to break this equation down 3x^2+30x+71
------------
It's not an equation, there's no equal sign.
I'll assume it's = 0.
----------------
I have no idea what "break this equation down" might mean.
---------------
The vertex is at the point where x = -b/2a
x = -30/6 = -5
f(-5) = 75 - 150 + 71 = -4
--> vertex at (-5,-4)